OpenStudy (anonymous):

Help please: Bob has two savings accounts. He deposited $200 more into account A than account B. After a period of time, account A has earned $84 in interest at 3%, and account B has earned $24 in interest at a rate of 2%. Write an equation to represent the situation. Explain each variable used. How much money did Bob initially deposit into each account? Solve the equation.

OpenStudy (anonymous):

If you could solve with steps that would be great.

OpenStudy (anonymous):

Could you help me please ? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is your variable?

OpenStudy (anonymous):

I didn"t pick one as yet. You could pick one if you would like.

jimthompson5910 (jim_thompson5910):

what's something that is unknown in this problem?

OpenStudy (anonymous):

The principle?

jimthompson5910 (jim_thompson5910):

for both accounts A and B so let x = amount deposited in account A y = amount deposited in account B

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

it says "He deposited $200 more into account A than account B" so that means x = y+200

jimthompson5910 (jim_thompson5910):

"After a period of time, account A has earned $84 in interest at 3%" so, I = P*r*t 84 = x*0.03*t 84 = 0.03*x*t

OpenStudy (anonymous):

ohhh okay

jimthompson5910 (jim_thompson5910):

also "account B has earned $24 in interest at a rate of 2%" I = P*r*t 24 = y*0.02*t 24 = (x+200)*0.02*t

jimthompson5910 (jim_thompson5910):

So your two equations are 84 = 0.03*x*t 24 = (x+200)*0.02*t

jimthompson5910 (jim_thompson5910):

you can use them both to solve for x and t

OpenStudy (anonymous):

okayy thanks so much! Now i understand what i did wrong when setting it up!

jimthompson5910 (jim_thompson5910):

you're welcome